OpenStudy (19rutlehop):

If a conditional statement is true, then its contrapositive __________. is true is false is both true and false

OpenStudy (prince1342):

This is tricky. I'm thinking D but not too sure

OpenStudy (littlet):

I think it is both true and false.

OpenStudy (ebonygotthejuice):

i think its false

OpenStudy (haseeb96):

i have no idea :D

OpenStudy (19rutlehop):

mf.

OpenStudy (19rutlehop):

It's false .

OpenStudy (dayakar):

CAN U GIVE ONE EXAMPLE

OpenStudy (19rutlehop):

^ you mad bro?

OpenStudy (dayakar):

conditional and contra positive are equivalent statements

OpenStudy (dayakar):

conditional: \[2>5\rightarrow 4 \times 2 =10 \left( TRUE \right)\] CONTRA POSITIVE: \[4 \times 2\neq 10 \rightarrow 2>5 \left( TRUE \right)\]

OpenStudy (19rutlehop):

k